# Evidence of Memory Effects in the Dynamics of Two-Level System Defect Ensembles Using Broadband, Cryogenic Transient Dielectric Spectroscopy

### v2

@@ Abstract
Two-level system (TLS) defects in dielectrics are a major source of decoherence in superconducting circuits, yet their atomistic origin, frequency distribution, and dipole moments remain poorly understood. Current probes, which are predominantly based on qubits or resonators, require complex fabrication and only measure defects within a narrow frequency band and limited mode volume, hindering direct insight into TLS defect behaviour in isolated materials and interfaces. Here, we introduce a broadband 3D waveguide spectroscopy technique that enables cryogenic probing of ensembles of TLS defects that we call Broadband Cryogenic Transient Dielectric Spectroscopy (BCTDS). Complementary to the dielectric dipper method, this approach probes a broader spectrum and reveals interference of drive-induced sidebands of the ensembles of TLS defects. The broadband and power-tunable nature of BCTDS makes it especially well-suited to the study of dressed-state physics in driven ensembles of TLS defects, including multi-photon processes and sideband-resolved dynamics. Additionally, BCTDS enables the identification of eigenmode frequencies of the undriven ensembles of TLS defects through characteristic V-shaped features obtained via Fourier analysis of time-domain signals, and shows evidence of memory effects arising from interactions and the broadband nature of our approach. Crucially, our method is modular and can be applied throughout the device fabrication process, informing mitigation strategies and advancing the design of low-loss materials with broad implications for quantum technologies and materials science.
@@ I Introduction
TLS defects are necessary to describe anomalous low-temperature properties of disordered solids, and they play a critical role in limiting the coherence of quantum technologies such as superconducting qubits and resonators. At low temperatures, many amorphous solids exhibit a thermodynamic behavior that deviates significantly from the predictions of Debye theory, which describes the specific heat of crystalline materials as scaling with . In contrast, experiments in the early 1970s revealed that glassy materials display a nearly linear temperature dependence of the specific heat at cryogenic temperatures [1]. To describe this low temperature behavior, Anderson, Halperin, and Varma [2], and independently Phillips [3], introduced the standard tunneling model, which treats disordered materials as a collection of effective double-well potentials where quantum tunneling can occur (Fig. 1a).
Today, it is widely believed that individual TLS defects are ubiquitous and form an ensemble whose collective low-energy excitations account for the observed linear temperature dependence of the specific heat capacity. Early models treated TLS defects as non-interacting and uniformly distributed in frequency, which sufficed to explain the specific heat capacity measurements. However, studies have shown that TLS defects can interact, leading to spectral diffusion, clustering, and nonlinear loss [5, 6, 7, 8, 9, 10, 11].
@@
While the notion of a TLS defect was initially developed to explain the thermodynamic properties of glassy materials, such defects have recently regained relevance in solid-state quantum computing technologies. This renewed interest is especially evident in superconducting circuits, which have become a leading platform for quantum computing primarily due to their fast gate times, design flexibility, and compatibility with advanced semiconductor fabrication techniques. However, their decoherence times are relatively short compared to other promising quantum technologies, such as neutral atoms, trapped ions, and spin qubits. From measurements of bulk loss tangents [12, 13], and modeling of the Purcell effect [14, 15, 16], it has become clear that other sources of loss must be present. Furthermore, measurements of qubit coherence while changing electric field and mechanical strain provide explicit evidence of TLS defects in superconducting circuits [17, 18, 19, 4, 20]. Today, it is widely believed that the dominant source of decoherence of superconducting circuits is TLS defects that reside on surfaces and interfaces with no clear consensus as to their specific atomistic origin [21, 22, 23, 24, 25] (Fig. 1b).
@@
TLS defects are typically probed via photonic or acoustic resonators [26, 27, 28, 29, 30] and qubits [12, 31, 17, 18, 19, 4]. These approaches provide high cooperativity but preferentially detect TLS defects with dipole moments aligned with the directly integrated probes, and are restricted to small bandwidths and mode volumes (Fig. 1c). Despite these limitations, several studies have demonstrated the existence of TLS-TLS interactions [32, 33, 19], and non-Markovian dynamics [34, 35], leading to questions about the nature of TLS defect interactions in the absence of strongly coupled probes [36]. In the spirit of the Mollow triplet, our setup enables the observation of dressed-state features in interacting ensembles of TLS defects, where the underlying spectral structure can reflect many-body dynamics rather than single-emitter coherence [37]. Furthermore, our method can be extended to accurately map the frequency and dipole moments of ensembles of TLS defects by employing a polarization-dependent drive. While conventional techniques such as electron spin resonance (ESR) focus on detecting net magnetization changes associated with spin-active defects using narrowband field excitation, BCTDS enables parallel, broadband interrogation of both spin and non-spin defects by directly measuring their transient electromagnetic and dielectric responses.
@@
II CRYOGENIC DIELECTRIC SPECTROSCOPY
Modular, Broadband 3D Microwave Waveguide
To probe the dielectric response in the microwave frequency domain, we leverage a broadband 3D waveguide (Fig. 2a). In contrast to qubit and resonator-based probes, this method allows for coupling to ensembles of TLS defects over a broad frequency range (Fig. 1c). As described in Appendix A, the closed waveguide comprises two WR-229 to coaxial cable adapters joined by a sample clamp. This modular approach allows us to easily measure samples of various sizes, shapes, and material compositions over a broadband frequency range of 3-6 GHz (Fig. 2b). Importantly, these features distinguish our approach from existing narrow-band qubit and resonator-based probes, which require samples to go through a full fabrication cycle, followed by wire bonding and packaging, making it nearly impossible to perform additional material processing and characterization. Furthermore, this constraint makes the development of new fabrication recipes wasteful and inefficient.
Dielectric Response of an Ensemble of TLS Defects
@@ II CRYOGENIC DIELECTRIC SPECTROSCOPY
To understand this system, we first consider a single TLS defect characterized by an energy detuning and a tunneling amplitude , as illustrated in Fig. 1a. The effective Hamiltonian for the -th TLS defect in the position basis can be written as [19]
@@
Cyrogenic Transient Dielectric Spectroscopy
In principle, BCTDS can be performed on any dielectric sample at any temperature. Furthermore, even metallic or superconducting samples with oxide layers can be probed in reflection (Appendix D, Fig. 9). Here, we will be primarily interested in the dielectric response of TLS defects, which solely arise at cryogenic temperatures, as described in Sec. I. Therefore, to extract information about the TLS defects, we perform BCTDS on our samples by inserting them into the waveguide (Fig. 2a) and cooling them to 10 mK using a dilution refrigerator. As mentioned in the previous section, driven dielectric spectroscopy monitors the response during the pulse, while transient dielectric spectroscopy monitors the post-pulse region (also known as the ring-down).
@@ III Results
Cryogenic Dielectric Spectroscopy of Different Materials
@@
As described in Sec. II, our waveguide enables a modular characterization of the driven and transient dielectric response of ensembles of TLS defects in various samples. The samples considered in this study started with a 50.8 mm diameter, 432 m thick, wafer of ultra-high-purity HEMEX sapphire from Crystal Systems, which has been measured to have bulk loss tangents of the order using the dielectric dipper technique [13]. We spin coat the samples with photoresist and dice them into rectangular chips with 27.5 5.5 mm dimensions. We clean the samples with solvents following a standard procedure described in Appendix B and mount them in the waveguide (Fig. 2a).
@@
Figure 3 summarizes the dielectric spectroscopy results obtained for different materials, alongside two room temperature control experiments of empty waveguide (Fig. 3a) and bare sapphire samples (Fig. 3b) that serves as a benchmark. Given that the TLS defects observed in our experiments have frequencies between 3 and 5 GHz, small temperature changes can significantly impact their thermal occupation, a key signature that has been explored in previous work [49, 50]. Furthermore, increases in temperature cause thermal occupation of phonon modes, which can lead to incoherent shaking of the double-well potential governing the TLS defects, likely broadening or destroying their spectral features. Our room temperature measurements (Fig. 3aii and bii) confirm this thermal saturation effect.
@@
As motivated in Sec. II, when we cool down to cryogenic temperatures and remove this thermal effect, linear response theory suggests that the spectral dips of the driven dielectric response should be corroborated by the longer-lived coherence times in the transient region. This is consistent with our findings when we cool down the same setups in Fig. 3a and b to 10 mK, we see an apparent narrowing of the spectral dips in the driven dielectric spectroscopy and a corresponding increase in the coherence of the transient dielectric response (Fig. 3c and d). These features are clear signatures of coherent energy storage and re-emission by the ensembles of TLS defects, which we discuss further in the context of master equation and Floquet theory in the next section (with further investigations in Appendix G). In Fig. 3c, we observe defect features in the absence of mounted samples, which we attribute to oxide layers present on the waveguide and antenna surfaces, and uncontrolled modes in the dilution fridge enclosure that couple to the waveguide. The signal in Fig. 3d likely originates from similar sources but may also include contributions from the solvent-cleaned sapphire samples in the waveguide, where residual surface contamination can result from leftover photoresist, adventitious carbon, or surface and dicing damages. In future work, we intend to remove these baseline contributions by retracting the samples in situ with a piezoelectric positioner, inspired by the dielectric dipper approach [13]. This calibration will enable further exploration of surface treatments such as etching, annealing, and passivation to reduce the measured transient dielectric response of bare sapphire and silicon wafers, which form the basis of all superconducting circuits [51, 52, 53].
@@
Today, most Josephson junctions are fabricated using various versions of double-angle evaporation, where the insulating Josephson junction layer is formed from a native oxide grown on deposited aluminum. However, all regions of the chips that contain aluminum will inadvertently grow such native oxides, and therefore, most superconducting circuits inevitably contain large areas of aluminum oxide. To emulate the native oxide in Josephson junctions and on aluminum metal layers, we grow a thin, 2 nm aluminum oxide layer in an atomic layer deposition tool on a bare sapphire wafer (Fig. 3b and d). The results of these dielectric spectroscopy measurements are shown in Fig. 3e. Here, we see clear signatures of ensembles of TLS defects, primarily in the transient region. The transient response exhibits pronounced collapse and revivals due to the interplay between external driving and TLS-defect ensemble interactions, clearly indicating memory effects [4, 54]. Furthermore, we want to highlight that the deposited 2 nm of aluminum oxide is approximately times the volume of the substrate, but is composed of the same atomic species. These findings demonstrate the importance of material morphology in the existence and behavior of TLS defects and motivate future work on passivating metallic layers [55] and developing crystalline insulating layers for Josephson junctions. This work also suggests why tantalum, which hosts a thin, stoichiometric native oxide, is an advantageous material choice for high-coherence superconducting circuits [51, 25].
@@
Finally, we also characterize the dielectric response of 1-4 m of photoresist on sapphire, as shown in Fig. 3f. This investigation is especially relevant because most resonators and waveguides leverage photolithography, meaning residual photoresist may be left on substrates if improperly removed. These photoresists are complex polymer chains that likely contain a high density of TLS defects [56, 57] (a detailed description of sample preparation is included in Appendix B). By understanding the dielectric properties of the resist layer, we can better assess its impact on junction performance and guide fabrication and device design optimizations [43]. Therefore, in Fig. 3f, we observe more transient response, consistent with our hypothesis that photoresist has many TLS defects.
@@
Dielectric Response under Different Drive Lengths
@@
In order to better examine the relationship between absorption in the driven dielectric response and the transient emission, we zoom into a narrower frequency range of 3.9-4.4 GHz and vary the pulse lengths in Fig. 4, featuring 20 ns (Fig. 4a), 50 ns (Fig. 4b), and 200 ns (Fig. 4c) pulses. Increasing the drive length improves the spectral resolution and reveals finer resonant features, which manifest themselves as sharper absorption dips during the pulse. After the drive is turned off, a similar effect is seen in the transient emission, where we observe sharper and more prolonged ringdowns. Without controlled access to the physical parameters of the TLS defects, it is encouraging to observe changes in the TLS dynamics as a result of different pulses. These findings support the feasibility of pulse-based coherent control of ensembles of TLS defects.
@@
To explore the connection between spectral dips and the surrounding emission features, we performed numerical simulations of two interacting driven spins. The full pulse evolution was treated using a master equation, while the periodic portion of the drive was analyzed separately using Floquet formalism, as detailed in Appendix G and illustrated in Fig. 12. During the pulse, the system is dressed by the external periodic drive, resulting in the emergence of sidebands. The corresponding Floquet quasi-energies show clear dips (Fig. 12b), and prolonged ring-downs within the vicinity of these dips, as seen in Fig. 12a. In Fig. 13, we qualitatively reproduce the sharpening of ring-downs for different pulse durations as observed in the experiment.
Further Investigation of Cryogenic Transient Dielectric Spectroscopy Results

### v3

@@ Abstract
Two-level system (TLS) defects in dielectrics cause decoherence in superconducting circuits, yet their origin, frequency distribution, and dipole moments remain poorly understood. Current probes, primarily based on qubits or resonators, require complex fabrication and measure defects only within narrow frequency bands and limited mode volumes, restricting insight into TLS behavior in isolated materials and interfaces. We introduce Broadband Cryogenic Transient Dielectric Spectroscopy (BCTDS), a broadband 3D waveguide technique that enables probing of TLS ensembles at cryogenic temperatures. Complementary to the dielectric dipper method, this approach probes a broader spectrum and reveals interference of drive-induced sidebands in TLS ensembles. The broadband, power-tunable nature of BCTDS makes it well suited for studying dressed-state physics in driven TLS ensembles, including multi-photon processes and sideband-resolved dynamics. By analyzing Fourier-transformed time-domain signals, BCTDS reveals eigen-mode frequencies of undriven TLS ensembles through characteristic V-shaped features and uncovers memory effects arising from interactions and broadband excitation. The modular method can be applied throughout device fabrication, informing mitigation strategies and advancing the design of low-loss materials with broad implications for quantum technologies and materials science.
@@ I Introduction
TLS defects are necessary to describe anomalous low-temperature properties of disordered solids, and they play a critical role in limiting the coherence of quantum technologies such as superconducting qubits and resonators. At low temperatures, many amorphous solids exhibit a thermodynamic behavior that deviates significantly from the predictions of Debye theory, which describes the specific heat of crystalline materials as scaling with . In contrast, experiments in the early 1970s revealed that glassy materials display a nearly linear temperature dependence of the specific heat at cryogenic temperatures [1]. To describe this low temperature behavior, Anderson, Halperin, and Varma [2], and independently Phillips [3], introduced the standard tunneling model, illustrated in Fig. 1(a). The model treats disordered materials as a collection of effective double-well potentials where quantum tunneling can occur. Today, it is widely believed that individual TLS defects are ubiquitous and form an ensemble whose collective low-energy excitations account for the observed linear temperature dependence of the specific heat capacity. Early models treated TLS defects as non-interacting and uniformly distributed in frequency, which sufficed to explain the specific heat capacity measurements. However, studies have shown that TLS defects can interact, leading to spectral diffusion, clustering, and nonlinear loss [4, 5, 6, 7, 8, 9, 10].
@@
While the notion of a TLS defect was initially developed to explain the thermodynamic properties of glassy materials, such defects have recently regained relevance in solid-state quantum computing technologies. This renewed interest is especially evident in superconducting circuits, which have become a leading platform for quantum computing primarily due to their fast gate times, design flexibility, and compatibility with advanced semiconductor fabrication techniques. However, their decoherence times are relatively short compared to other promising quantum technologies, such as neutral atoms, trapped ions, and spin qubits. From measurements of bulk loss tangents [12, 13], and modeling of the Purcell effect [14, 15, 16], it has become clear that other sources of loss must be present. Furthermore, measurements of qubit coherence while changing electric field and mechanical strain provide explicit evidence of TLS defects in superconducting circuits [17, 18, 19, 11, 20]. Currently, the prevailing understanding is that decoherence in superconducting circuits is dominated by TLS defects on surfaces and interfaces, as illustrated in Fig. 1(b), although their specific atomistic origin remains unclear [21, 22, 23, 24, 25].
@@
TLS defects are typically probed via photonic or acoustic resonators [26, 27, 28, 29, 30] and qubits [12, 31, 17, 18, 19, 11]. These approaches provide high cooperativity but preferentially detect TLS defects with dipole moments aligned with the directly integrated probes, and are restricted to small bandwidths and mode volumes, as depicted in Fig. 1(c). Despite these limitations, several studies have demonstrated the existence of TLS-TLS interactions [32, 33, 19], and non-Markovian dynamics [34, 35], leading to questions about the nature of TLS defect interactions in the absence of strongly coupled probes [36]. In the spirit of the Mollow triplet, our setup enables the observation of dressed-state features in interacting ensembles of TLS defects, where the underlying spectral structure can reflect many-body dynamics rather than single-emitter coherence [37]. Furthermore, our method can be extended to accurately map the frequency and dipole moments of ensembles of TLS defects by employing a polarization-dependent drive. While conventional techniques such as electron spin resonance (ESR) focus on detecting net magnetization changes associated with spin-active defects using narrowband field excitation, BCTDS enables parallel, broadband interrogation of both spin and non-spin defects by directly measuring their transient electromagnetic and dielectric responses.
@@
II BROADBAND CRYOGENIC TRANSIENT DIELECTRIC SPECTROSCOPY
II.1 Modular, Broadband 3D Microwave Waveguide
To probe the dielectric response in the microwave frequency domain, we leverage a broadband 3D waveguide, shown in Fig. 2(a). In contrast to qubit and resonator-based probes, this method allows for coupling to ensembles of TLS defects over a broad frequency range. As described in Appendix A, the closed waveguide comprises two WR-229 to coaxial cable adapters joined by a sample clamp. This modular approach allows us to easily measure samples of different sizes, shapes, and material compositions across a broadband range of 3–6 GHz. The corresponding S-parameter curves are shown in Fig. 2(b). Importantly, these features distinguish our approach from existing narrow-band qubit and resonator-based probes, which require samples to go through a full fabrication cycle, followed by wire bonding and packaging, making it nearly impossible to perform additional material processing and characterization. Furthermore, this constraint makes the development of new fabrication recipes wasteful and inefficient.
II.2 Dielectric Response of an Ensemble of TLS Defects
@@
To understand this system, we first consider a single TLS defect characterized by an energy detuning and a tunneling amplitude , as illustrated in Fig. 1(a). The effective Hamiltonian for the -th TLS defect in the position basis can be written as [19]
@@
II.3 Experimental Implementation
In principle, BCTDS can be performed on any dielectric sample at any temperature. Furthermore, even metallic or superconducting samples with oxide layers can be probed in reflection (Appendix D, Fig. 9). Here, we will be primarily interested in the dielectric response of TLS defects, which solely arise at cryogenic temperatures, as described in Sec. I. Therefore, to extract information about the TLS defects, we perform BCTDS on our samples by inserting them into the waveguide and cooling them to 10 mK using a dilution refrigerator. As mentioned in the previous section, driven dielectric spectroscopy monitors the response during the pulse, while transient dielectric spectroscopy monitors the post-pulse region (also known as the ring-down).
@@ III Results
III.1 Cryogenic Dielectric Spectroscopy of Different Materials
@@
As described in Sec. II.1, our waveguide enables a modular characterization of the driven and transient dielectric response of ensembles of TLS defects in various samples. The samples considered in this study started with a 50.8 mm diameter, 432 m thick, wafer of ultra-high-purity HEMEX sapphire from Crystal Systems, which has been measured to have bulk loss tangents of the order using the dielectric dipper technique [13]. We spin coat the samples with photoresist and dice them into rectangular chips with 27.5 5.5 mm dimensions. We clean the samples with solvents following a standard procedure described in Appendix B and mount them in the waveguide.
@@
Figure 3 summarizes the dielectric spectroscopy results obtained for different materials. We perform room-temperature control experiments on an empty waveguide and on a waveguide with bare sapphire samples that serve as benchmarks, shown in Fig. 3(a) and 3(b), respectively. Given that the TLS defects observed in our experiments have frequencies between 3 and 5 GHz, small temperature changes can significantly impact their thermal occupation, a key signature that has been explored in previous work [49, 50]. Furthermore, increases in temperature cause thermal occupation of phonon modes, which can lead to incoherent shaking of the double-well potential governing the TLS defects, likely broadening or destroying their spectral features. Our room temperature measurements in Fig. 3(aii) and 3(bii) confirm this thermal saturation effect.
@@
As motivated in Sec. II.3, when we cool down to cryogenic temperatures and remove this thermal effect, linear response theory suggests that the spectral dips of the driven dielectric response should be corroborated by the longer-lived coherence times in the transient region. This is consistent with our findings when we cool down the same setups in Fig. 3(a) and 3(b) to 10 mK, we see an apparent narrowing of the spectral dips in the driven dielectric spectroscopy and a corresponding increase in the coherence of the transient dielectric response, shown in Fig. 3(c) and 3(d). These features are clear signatures of coherent energy storage and re-emission by the ensembles of TLS defects, which we discuss further in the context of master equation and Floquet theory in the next section (and with further investigations in Appendix G). In Fig. 3(c), we observe defect features in the absence of mounted samples, which we attribute to oxide layers present on the waveguide and antenna surfaces, and uncontrolled modes in the dilution fridge enclosure that couple to the waveguide. The signal in Fig. 3(d) likely originates from similar sources but may also include contributions from the solvent-cleaned sapphire samples in the waveguide, where residual surface contamination can result from leftover photoresist, adventitious carbon, or surface and dicing damages. In future work, we intend to remove these baseline contributions by retracting the samples in situ with a piezoelectric positioner, inspired by the dielectric dipper approach [13]. This calibration will enable further exploration of surface treatments such as etching, annealing, and passivation to reduce the measured transient dielectric response of bare sapphire and silicon wafers, which form the basis of all superconducting circuits [51, 52, 53].
@@
Today, most Josephson junctions are fabricated using various versions of double-angle evaporation, where the insulating Josephson junction layer is formed from a native oxide grown on deposited aluminum. However, all regions of the chips that contain aluminum will inadvertently grow such native oxides, and therefore, most superconducting circuits inevitably contain large areas of aluminum oxide. To emulate the native oxide in Josephson junctions and on aluminum metal layers, we grow a thin, 2 nm aluminum oxide layer in an atomic layer deposition tool on a bare sapphire wafer. The results of these dielectric spectroscopy measurements are shown in Fig. 3(e). Here, we see clear signatures of ensembles of TLS defects, primarily in the transient region. The transient response exhibits pronounced collapse and revivals due to the interplay between external driving and TLS-defect ensemble interactions, clearly indicating memory effects [11, 54]. Furthermore, we want to highlight that the deposited 2 nm of aluminum oxide is approximately times the volume of the substrate, but is composed of the same atomic species. These findings demonstrate the importance of material morphology in the existence and behavior of TLS defects and motivate future work on passivating metallic layers [55] and developing crystalline insulating layers for Josephson junctions. This work also suggests why tantalum, which hosts a thin, stoichiometric native oxide, is an advantageous material choice for high-coherence superconducting circuits [51, 25].
@@
Finally, we also characterize the dielectric response of 1-4 m of photoresist on sapphire, as shown in Fig. 3(f). This investigation is especially relevant because most resonators and waveguides leverage photolithography, meaning residual photoresist may be left on substrates if improperly removed. These photoresists are complex polymer chains that likely contain a high density of TLS defects [56, 57] (a detailed description of sample preparation is included in Appendix B). By understanding the dielectric properties of the resist layer, we can better assess its impact on junction performance and guide fabrication and device design optimizations [43]. Therefore, in Fig. 3(f), we observe more transient response, consistent with our hypothesis that photoresist has many TLS defects.
@@
III.2 Dielectric Response under Different Drive Lengths
@@
In order to better examine the relationship between absorption in the driven dielectric response and the transient emission, we zoom into a narrower frequency range of 3.9-4.4 GHz and vary the pulse lengths in Fig. 4(a)-4(c), featuring 20 ns, 50 ns, and 200 ns pulses. Increasing the drive length improves the spectral resolution and reveals finer resonant features, which manifest themselves as sharper absorption dips during the pulse. After the drive is turned off, a similar effect is seen in the transient emission, where we observe sharper and more prolonged ringdowns. Without controlled access to the physical parameters of the TLS defects, it is encouraging to observe changes in the TLS dynamics as a result of different pulses. These findings support the feasibility of pulse-based coherent control of ensembles of TLS defects.
@@
To explore the connection between spectral dips and the surrounding emission features, we performed numerical simulations of two interacting driven spins. The full pulse evolution was treated using a master equation, while the periodic portion of the drive was analyzed separately using Floquet formalism, as detailed in Appendix G and illustrated in Fig. 12. During the pulse, the system is dressed by the external periodic drive, resulting in the emergence of sidebands. The corresponding Floquet quasi-energies show clear dips in Fig. 12(b), and prolonged ring-downs within the vicinity of these dips, as seen in Fig. 12(a). In Fig. 13, we qualitatively reproduce the sharpening of ring-downs for different pulse durations as observed in the experiment.
III.3 Further Investigation of Cryogenic Transient Dielectric Spectroscopy Results